OpenStudy (anonymous):

This morning, Elena wrote a check in the amount of $42.50. This brought her balance down to –$14.25. Suppose n represents Elena's balance before she wrote the check?? Which equation can be used to represent this situation? A. n – 14.25 = 42.50 B. n + 42.50 = –14.25 C. n – 42.50 = –14.25 D. n – 14.25 = –42.50

OpenStudy (anonymous):

i think its b honestly.

OpenStudy (sdfgsdfgs):

But writing a check means taking money out which means the amount sho6uld be add o5r subtract from the total n?

OpenStudy (anonymous):

true

OpenStudy (sdfgsdfgs):

So which u think is the right ans now?

OpenStudy (holly00d1248):

think of it this way your starting value N minus a number 42.50 equals a number 14.25 does that help

OpenStudy (anonymous):

It Is C

OpenStudy (anonymous):

yes thx im working on it now brb..

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

you guys are awsome

OpenStudy (holly00d1248):

your welcome

OpenStudy (anonymous):

i have a notherone

OpenStudy (anonymous):

i dont understand it ;( Use transformations to solve the equation. c – 16 = 41 A. c = –57 B. c = –25 C. c = 25 D. c = 57

OpenStudy (sdfgsdfgs):

add 16 on the left and right hand side of the eqn n what will u get?

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

well you should get 0 because it is a negative

OpenStudy (sdfgsdfgs):

c – 16 = 41 then adding 16 on the left n right c - 16 + 16 = 41 + 16 what will u get now?

OpenStudy (anonymous):

;yes

OpenStudy (anonymous):

thx

OpenStudy (sdfgsdfgs):

So which is the right ans?

OpenStudy (anonymous):

hold on im working it out im close

OpenStudy (anonymous):

D

OpenStudy (anonymous):

thats what i think it is tha ans

OpenStudy (sdfgsdfgs):

That is correct. Good job! :)

OpenStudy (anonymous):

Thx

OpenStudy (anonymous):

another?

OpenStudy (anonymous):

What is the solution to the equation? a + (–0.4) = –3 a =

OpenStudy (sdfgsdfgs):

Try this: a + (–0.4) = –3 a + (–0.4) - (-0.4) = –3 - (-0.4)

OpenStudy (anonymous):

-2.6

OpenStudy (anonymous):

u are the best @MariahDimond and Smart

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

A=-2.4

OpenStudy (anonymous):

Use a transformation to solve the equation. x – 0.7 = 1.9 A. x = –2.6 B. x = –1.2 C. x = 1.2 D. x = 2.6

OpenStudy (sdfgsdfgs):

+ 0.7 on left and right. What will u get?

OpenStudy (anonymous):

0.7+0.7

OpenStudy (anonymous):

x – 0.7+0.7= 1.9+0.7

OpenStudy (sdfgsdfgs):

x – 0.7 = 1.9 x – 0.7 + 0.7 = 1.9 + 0.7 @MariahDimond yes :) good job!

OpenStudy (anonymous):

:) THX

OpenStudy (anonymous):

X=1.2

OpenStudy (anonymous):

What is the solution to the equation? y – 0.7 = –2.2 y =

OpenStudy (anonymous):

y-0.7+0.7=-2.2 0.7

OpenStudy (anonymous):

Y=2.9

OpenStudy (sdfgsdfgs):

U are getting the hang of it now @MariahDimond

OpenStudy (anonymous):

@sdfgsdfgs Yes i am

OpenStudy (anonymous):

Use a transformation to solve the equation. s – 14.2 = –18.7 A. s = –32.9 B. s = –4.5 C. s = 4.5 D. s = 32.9

OpenStudy (anonymous):

s – 14.2+14.2 = –18.7 +14.2

OpenStudy (sdfgsdfgs):

Looks good - so what is the right ans?

OpenStudy (anonymous):

S=32.9

OpenStudy (sdfgsdfgs):

Nope....u forgot there is a - sign in front of 18.7

OpenStudy (anonymous):

@sdfgsdfgs now i have english

OpenStudy (sdfgsdfgs):

Wait plz re-do the last one...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Use a transformation to solve the equation. s – 14.2 = –18.7 A. s = –32.9 B. s = –4.5 C. s = 4.5 D. s = 32.9

OpenStudy (sdfgsdfgs):

u forgot there is a - sign in front of 18.7

OpenStudy (anonymous):

s – 14.2+14.2 =–18.7+14.2

OpenStudy (anonymous):

a negative?

OpenStudy (sdfgsdfgs):

dont worry. keep going plz :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

S=32.9

OpenStudy (anonymous):

i gtg

OpenStudy (sdfgsdfgs):

It is not right so plz post the last Q again when u are back @MariahDimond

OpenStudy (anonymous):

@sdfgsdfgs ok

OpenStudy (anonymous):

wait S=-32.9